# On the Complexity of Bipartite Degree Realizability

### v1

@@ Abstract
We further show that, assuming the NP-completeness of unrestricted BDR, the problem $\mathrm{BDR}_{c_{1},c_{2}}$ remains NP-complete for all $0<c_{2}<\tfrac{1}{2}$ and $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$. Our results clarify the algorithmic landscape of bipartite degree realization and contribute to the broader study of potentially bipartite graphic degree sequences.
@@ 1 Introduction
A nonnegative integer sequence $D=(d_{1},\dots,d_{n})$ is called graphic if there exists a simple graph on $n$ vertices whose degrees are exactly $d_{i}$. We call such a graph a realization of $D$. Similarly, a pair of degree sequences $(D_{1},D_{2})$ is graphic if it has a bipartite graph realization, $G=(U,V,E)$, such that the degrees in $U$ are $D_{1}$ and the degrees in $V$ are $D_{2}$. In 1981, Rao considered the potentially and forcibly $\mathcal{P}$-graphic degree sequences [9]. Given a graph property $\mathcal{P}$ and a degree sequence $D$, we say that $D$ is potentially $\mathcal{P}$-graphic if $D$ has a simple graph realization with property $\mathcal{P}$. Similarly, $D$ is forcibly $\mathcal{P}$-graphic if any realization of $D$ has property $\mathcal{P}$. For example, if the sum of the degrees in a degree sequence $D$ with length $n$ is less than 2n22n-2, then $D$ is forcibly disconnected. If the sum is at least 2n22n-2, and there is no 0 degree, then $D$ is potentially connected. If the minimum degree is greater than $\left\lfloor\frac{n-2}{2}\right\rfloor$ then $D$ is forcibly connected.
One of the open problems that Rao asked was the potentially bipartite degree sequences. According to his terminology, we call a degree sequence potentially bipartite if it admits at least one bipartite realization, and forcibly not bipartite if none of its realizations is bipartite. We will denote the potentially bipartite degree sequence problem as the BDR problem, that is, the Bipartite Degree Realization problem. The input of the BDR problem is a degree sequence $D$, and the problem is to decide if $D$ has a bipartite graph realization. This problem has an unknown computational complexity till today, although the following results are well-known. The Gale-Ryser theorem provides a polynomial time algorithm to decide if a pair of degree sequences $(D_{1},D_{2})$ has a bipartite graph realization $G=(U,V,E)$, such that the degrees in $U$ are $D_{1}$ and the degrees in $V$ are $D_{2}$ [5, 10]. Also, there is a polynomial running time dynamic programming algorithm to decide if a degree sequence $D$ can be split into two degree sequences, $D_{1}$ and $D_{2}$ with equal sum, a necessary condition to have a bipartite graph realization. However, there might be an exponential number of solutions how to split $D$ into two degree sequences with equal sums [2], and it is not trivial how to decide if any of them satisfies the sufficient and necessary conditions given by Gale and Ryser.
@@
We also show the following. If BDR is NP-complete, then any parametric version $\textsc{BDR}_{c_{1},c_{2}}$ with $c_{1}\leq\frac{1}{2}$ and $c_{2}>\sqrt{2c_{1}}-c_{1}$ is NP-complete as well. This result follows from a simple polynomial-time padding (or scaling) reduction that preserves bipartite realizability.
@@
The computational complexity of deciding whether a degree sequence admits a bipartite realization has recently been studied also by Bar-Noy et al. in a general framework that includes both simple graphs and multigraphs [2]. Our work is concerned exclusively with simple graphs. Accordingly, we restrict attention to those results of Bar-Noy et al. that apply to simple bipartite graphs and are directly comparable to ours. In this restricted setting, Bar-Noy et al. identify polynomial-time solvable cases primarily under strong structural assumptions on the degree sequence, most notably when the sequence is nearly regular or exhibits only small deviations from its average degree. In particular, their tractable cases for simple bipartite graphs rely on bounded variance conditions or on the existence of a near-balanced partition in which the degrees on each side differ by at most a constant or sublinear amount. These results emphasize global balance and regularity rather than extremal degree constraints. In contrast, the present work establishes polynomial-time decidability for bipartite realizability under explicit extremal degree conditions of the form $c_{1}n\leq d_{i}\leq c_{2}n$, for constants $0\leq c_{1}\leq c_{2}<1$, covering a broad class of dense degree sequences with linear minimum and maximum degree bounds. Our results apply even when the degree sequence exhibits large variance, provided all degrees remain within the prescribed interval. This places our tractability region in a fundamentally different part of the parameter space than that considered by Bar-Noy et al. for simple graphs.
@@
In another work, Bar-Noy et al. also proved that the corresponding bipartite realization problem for multigraphs is NP-complete [3]. In that setting, degrees are allowed to be arbitrarily large and are not bounded as a function of the number of vertices. From this perspective, the hardness result aligns with the inherent difficulty of partitioning sequences of large integers into two parts with equal sums, a problem that is already NP-complete [6]. Consequently, these multigraph results are not directly comparable to the simple graph setting considered here.
@@
Bar-Noy et al. leave open the computational complexity of bipartite realizability for simple graphs when degrees are bounded only by n1n-1 [3]. Our results contribute to this open direction by showing that if bipartite realizability for simple graphs with degrees bounded by n1n-1 were NP-complete, then NP-completeness would persist even under much stronger restrictions, namely for any upper bound of the form $cn$ for any fixed constant $c>0$.
@@ 2 Preliminaries
We introduce the necessary ingredients to achieve the stated results. We are going to change graphs using hinge-flips. A hinge-flip operation removes an edge $(u_{1},v)$ and add an edge $(u_{2},v)$ not presented in the edge set before the operation. The hinge-flip operations do not change the sum of the degrees, however, they do change the degree sequence of the graph on which they act. However, observe the following: if in a pair of degree sequences $(D_{1},D_{2})$ the degree of $u_{i}$ is greater than the degree of $u_{j}$, then in any realization of $(D_{1},D_{2})$ there exists a hinge-flip that removes an edge $(u_{i},v)$ and adds an edge $(u_{j},v)$. That is, if a pair of degree sequences $(D_{1},D_{2})$ is graphic than any pair of degree sequences $(D_{1}^{\prime},D_{2})$ is also graphic that can be obtained from $(D_{1},D_{2})$ by decreasing a degree $d_{j}\in D_{1}$ and increasing a $d_{i}\in D_{1}$ with $d_{i}<d_{j}$. When $d_{i}<d_{j}-1$, we call the corresponding hinge-flip a balancing hinge-flip (if $d_{i}=d_{j}+1$, the hinge flip only swap the two degrees). Observe that a balancing hinge-flip does not change the sum of the degrees, cannot decrease the minimum degree as well as cannot increase the maximum degree. This observation leads to the notation of the least balanced degree sequence and a lemma on it.
Definition 1.
@@
Lemma 1.
Let $(D_{1},D_{2})$ be a pair of degree sequences. If $D_{1}\in\mathcal{D}(n_{1},\Sigma,\Delta_{1},d_{1})$, $D_{2}\in\mathcal{D}(n_{2},\Sigma,\Delta_{2},d_{2})$ and $(LBDS(n_{1},\Sigma,\Delta_{1},d_{1}),LBDS(n_{2},\Sigma,\Delta_{2},d_{2}))$ is graphic then $(D_{1},D_{2})$ is also graphic.
The proof can be found in [8]; an analogous observation on simple degree sequences was already observed by Aigner and Triesch [1].
The graphicality of a bipartite degree sequence can be checked with the Gale-Ryser theorem [5, 10]. The Gale-Ryser theorem says that a bipartite degree sequence $D:=(D_{1},D_{2})$ with $D_{1}=d_{1,1}\geq d_{1,2}\geq\ldots\geq d_{1,n}$ and $D_{2}=d_{2,1},d_{2,2},\ldots,d_{2,m}$ is graphic if and only if the sum of the degrees in $D_{1}$ equals the sum of the degrees in $D_{2}$ and for all $k=1,2,\ldots,n$,
$\sum_{i=1}^{k}d_{1,i}\leq\sum_{j=1}^{m}\min\{k,d_{2,j}\}.$ (1)
We make the following observation. A bipartite degree sequence $D:=(D_{1},D_{2})$ with $D_{1}=d_{1,1}\geq d_{1,2}\geq\ldots\geq d_{1,n}$ and $D_{2}=d_{2,1},d_{2,2},\ldots,d_{2,m}$ is graphic if and only if the degree sequence $D^{\prime}:=d_{1,1}+n-1,d_{1,2}+n-1,\ldots,d_{1,n}+n-1,d_{2,1},d_{2,2},\ldots,d_{2,m}:=$ $d^{\prime}_{1},d^{\prime}_{2},\ldots d^{\prime}_{n^{\prime}}$ is graphic. Then for all $k\leq n$, the Gale-Ryser inequalities in equation 1 are equivalent with the Erdős-Gallai inequalities [4]:
$\sum_{i=1}^{k}d^{\prime}_{i}\leq k(k-1)+\sum_{j=k+1}^{n^{\prime}}\min\{k,d_{j}\}.$ (2)
@@
Tripathi and Vijay [11] proved that a degree sequence is graphic if and only if the sum of the degrees is even and for all $k$ with $d_{k}>d_{k+1}$ the Erdős-Gallai inequalities hold. It leads to the following observation.
Observation 1.
The bipartite degree sequence $D=(LDBS(n_{1},\Sigma,\Delta_{1},d_{1}),LDBS(n_{2},\Sigma,\Delta_{2},d_{2}))$ is graphic if and only if the Gale-Ryser inequality with $k=\left\lfloor\frac{\Sigma-n_{1}d_{1}}{\Delta_{1}-d_{1}}\right\rfloor$ and $k=\left\lceil\frac{\Sigma-n_{1}d_{1}}{\Delta_{1}-d_{1}}\right\rceil$ are satisfied.
@@ 3 Polynomial algorithm for $\textsc{BDR}_{c_{1},c_{2}}$
We are going to prove the tractable result of the parameterized BDR problem.
Theorem 2.
@@
Observe the following. The inequality $c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$ implies that $c_{1}\leq\frac{1}{2}$ because for any $c_{1}>\frac{1}{2}$ it holds that $c_{1}>\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$. When $0\leq c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$ , it is enough to show that for any $s_{1}\leq s_{2}$ with $s_{1}+s_{2}=n$, and for any $\Sigma$ with $c_{1}s_{2}\leq\Sigma\leq c_{2}s_{1}$, the bipartite degree sequence $D:=(LBDS(s_{1},\Sigma,c_{1}n,c_{2}n),LBDS(s_{2},\Sigma,c_{1}n,c_{2}n))$ is graphic. Indeed, according to Lemma1, then any degree sequence $D^{\prime}:=(D_{1},D_{2})$ with $D_{1}\in\mathcal{D}(s_{1},\Sigma,c_{1}n,c_{2}n)$ and $D_{2}\in\mathcal{D}(s_{2},\Sigma,c_{1}n,c_{2}n)$ is graphic. Therefore, a degree sequence $D\in\mathcal{D}(n,2\Sigma,c_{1}n,c_{2}n)$ is graphic if and only if $D$ has a split of two degree sequences with equal sum. This is the subset sum problem, which is NP-complete in general, but has a pseudo-polynomial time algorithm. The pseudopolinomiality means that the runtime of the algorithm grows polynomial with the values of the numbers and not the number of digits needed to describe the numbers. However, any degree is at most $\frac{n}{2}$, therefore the subset sum problem is actually in P for these degree sequences.
We use the Gale-Ryser theorem to prove that the bipartite degree sequence $D$ is graphic. According to Observation 1, it is enough to check the Gale-Ryser inequalities with $k=\left\lfloor\frac{\Sigma-s_{1}c_{1}n}{(c_{2}-c_{1})n}\right\rfloor$ and $k=\left\lceil\frac{\Sigma-s_{1}c_{1}n}{(c_{2}-c_{1})}\right\rceil$.
@@
$c_{1}\geq\frac{c_{2}^{2}}{1-c_{1}},$
which is equivalent with $c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$. There are possible intermediate degrees with
@@
We have to check those Gale-Ryser inequalities which were obtained in Observation 1. The first is:
$n_{1}c_{2}n\;\leq\;n_{2}\min\{n_{1},c_{2}n\}\;+\;\min\{n_{1},d_{2}\}\;+\;m_{2}\min\{n_{1},c_{1}n\}.$ ($\star$)
We now prove $(\star)$ by splitting into three exhaustive cases according to the relative size of $n_{1}$.
@@
In this range we have $\min\{n_{1},c_{2}n\}=c_{2}n$, $\min\{n_{1},d_{2}\}=d_{2}$ (because $d_{2}<c_{2}n<n_{1}$) and $\min\{n_{1},c_{1}n\}=c_{1}n$. Thus, the right-hand side of $(\star)$ equals
@@
This is exactly the desired inequality in Case (A). Hence $(\star)$ holds when $n_{1}\geq c_{2}n$.
@@
Here $\min\{n_{1},c_{2}n\}=n_{1}$, $\min\{n_{1},d_{2}\}=n_{1}$ (since $d_{2}\geq c_{1}n\geq n_{1}$), and $\min\{n_{1},c_{1}n\}=n_{1}$. Therefore the right-hand side of $(\star)$ equals
@@
However, we know that $c_{2}\leq\frac{c_{1}}{c_{2}+c_{1}}$ for it is equivalent with the condition $\frac{c_{2}^{2}}{1-c_{2}}\leq c_{1}$. Therefore, $(\star)$ holds in Case (B).
@@
In this range we have $\min\{n_{1},c_{2}n\}=n_{1}$, and $\min\{n_{1},c_{1}n\}=c_{1}n$. Hence $(\star)$ becomes
@@
$n_{1}c_{2}n+d_{1}+\left(\frac{n}{2}-n_{1}+x-1\right)c_{1}n=(n_{1}+y)c_{2}n+d_{2}+\left(\frac{n}{2}-n_{1}-x-y-1\right)c_{1}n.$ (5)
@@
$n_{1}c_{2}n\leq\left(n_{1}+\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}\right)n_{1}+\min\{n_{1},d_{2}\}+\left(\frac{n}{2}-n_{1}-x-\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}-1\right)c_{1}n.$
@@
$0\leq\underbrace{\eta^{2}-(c_{1}+c_{2})\eta+\frac{c_{1}}{2}}_{R_{1}}+\underbrace{\frac{\frac{d_{1}-d_{2}}{n^{2}}}{c_{2}-c_{1}}(\eta-c_{1})+\frac{\min\{\eta,\frac{d_{2}}{n}\}}{n}}_{R_{2}}+\underbrace{z\left(\frac{2c_{1}}{c_{2}-c_{1}}\eta-\left(1+\frac{2c_{1}}{c_{2}-c_{1}}\right)c_{1}\right)}_{R_{3}}.$ (6)
@@
Therefore $R_{2}$ is always positive. Therefore, we can look at equation 6 as

### v2

@@ Abstract
We further show that, assuming the NP-completeness of unrestricted BDR, the problem $\mathrm{BDR}_{c_{1},c_{2}}$ remains NP-complete for all $0<c_{2}<\frac{1}{2}$ and $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$. Our results clarify the algorithmic landscape of bipartite degree realization and contribute to the broader study of potentially bipartite graphic degree sequences.
keywords:
graph realizations, bipartite graphs, degree sequences, bigraphic sequences, parameterized complexity
{MSCcodes}
05C07, 05C85, 68Q25
@@ 1 Introduction
A nonnegative integer sequence $D=(d_{1},\dots,d_{n})$ is called graphic if there exists a simple graph on $n$ vertices whose degrees are exactly $d_{i}$. We call such a graph a realization of $D$. Similarly, a pair of degree sequences $(D_{1},D_{2})$ is bigraphic if it has a bipartite graph realization, $G=(U,V,E)$, such that the degrees in $U$ are $D_{1}$ and the degrees in $V$ are $D_{2}$. In 1981, Rao considered the potentially and forcibly $\mathcal{P}$-graphic degree sequences [rao]. Given a graph property $\mathcal{P}$ and a degree sequence $D$, we say that $D$ is potentially $\mathcal{P}$-graphic if $D$ has a simple graph realization with property $\mathcal{P}$. Similarly, $D$ is forcibly $\mathcal{P}$-graphic if any realization of $D$ has property $\mathcal{P}$. For example, if the sum of the degrees in a degree sequence $D$ with length $n$ is less than 2n22n-2, then $D$ is forcibly disconnected. If the sum is at least 2n22n-2, and there is no 0 degree, then $D$ is potentially connected. If the minimum degree is greater than $\left\lfloor\frac{n-2}{2}\right\rfloor$ then $D$ is forcibly connected.
One of the open problems that Rao asked was the potentially bipartite degree sequences. According to his terminology, we call a degree sequence potentially bipartite if it admits at least one bipartite realization, and forcibly not bipartite if none of its realizations is bipartite. We will denote the potentially bipartite degree sequence problem as the BDR problem, that is, the Bipartite Degree Realization problem. The input of the BDR problem is a degree sequence $D$, and the problem is to decide if $D$ has a bipartite graph realization. This problem has an unknown computational complexity till today, although the following results are well-known. The Gale-Ryser theorem provides a polynomial time algorithm to decide if a pair of degree sequences $(D_{1},D_{2})$ has a bipartite graph realization $G=(U,V,E)$, such that the degrees in $U$ are $D_{1}$ and the degrees in $V$ are $D_{2}$ [gale, ryser]. Also, there is a polynomial running time dynamic programming algorithm to decide if a degree sequence $D$ can be split into two degree sequences, $D_{1}$ and $D_{2}$ with equal sum, a necessary condition to have a bipartite graph realization. However, there might be an exponential number of solutions how to split $D$ into two degree sequences with equal sums [barnoy1], and it is not trivial how to decide if any of them satisfies the sufficient and necessary conditions given by Gale and Ryser.
@@
We also show the following. If BDR is NP-complete, then any parametric version $\textsc{BDR}_{c_{1},c_{2}}$ with $c_{1}\leq\frac{1}{2}$ and $c_{2}>\sqrt{2c_{1}}-c_{1}$ is NP-complete as well (we will show that $c_{2}>\sqrt{2c_{1}}-c_{1}$ iff $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$). This result follows from a simple polynomial-time padding (or scaling) reduction that preserves bipartite realizability.
@@
The computational complexity of deciding whether a degree sequence admits a bipartite realization has recently been studied also by Bar-Noy et al. in a general framework that includes both simple graphs and multigraphs [barnoy1]. Our work is concerned exclusively with simple graphs. Accordingly, we restrict attention to those results of Bar-Noy et al. that apply to simple bipartite graphs and are directly comparable to ours. In this restricted setting, Bar-Noy et al. identify polynomial-time solvable cases primarily under strong structural assumptions on the degree sequence, most notably when the sequence has a high-low partition. A high-low partition of a degree sequence $D$ is a partition into $D_{1}$ and $D_{2}$ such that for all $d_{i}\in D_{1}$ and $d_{j}\in D_{2}$, $d_{i}\leq d_{j}$. A high-low partiton $(D_{1},D_{2})$ is well-behaved if the maximum degree in each vertex class is at most the size of the other vertex class. The main result of Bar-Noy et al. [barnoy1] is the following. If $D$ has a well-behaved high-low partition, then $D$ has a bipartite graph realization if and only if the well-behaved high-low partition has a bipartite graph realization. That is, if the well-behaved high-low partition does not have a bipartite graph realization then none of the equal sum partitions of the underlying degree sequence is bigraphic. These results assume a special global property (that is, well-behaved high-low partition) rather than extremal degree constraints.
In contrast, in the present work, we identify degree sequence classes in which any partition into equal sums has a bipartite graph realization. This establishes polynomial-time decidability for bipartite realizability under explicit extremal degree conditions of the form $c_{1}n\leq d_{i}\leq c_{2}n$, for constants $0\leq c_{1}\leq c_{2}<1$, covering a broad class of dense degree sequences with linear minimum and maximum degree bounds. Our results apply when a degree sequence does not have a high-low partition, and also, even when the degree sequence exhibits large variance, provided all degrees remain within the prescribed interval. The largest difference between the minimum and maximum degrees could be obtained by maximizing $\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}n-c_{1}n$ which turns out to be $(3-2\sqrt{2})n\approx 0.17157n$ when $c_{1}=\frac{3\sqrt{2}}{2}-2$. This places our tractability region in a fundamentally different part of the parameter space than that considered by Bar-Noy et al. for simple graphs.
@@
In another work, Bar-Noy et al. also proved that the corresponding bipartite realization problem for multigraphs is NP-complete [barnoy2]. In that setting, degrees are allowed to be arbitrarily large and are not bounded as a function of the number of vertices. From this perspective, the hardness result aligns with the inherent difficulty of partitioning sequences of large integers into two parts with equal sums, a problem that is already NP-complete [garey-johnson]. Consequently, these multigraph results are not directly comparable to the simple graph setting considered here.
@@
Bar-Noy et al. leave open the computational complexity of bipartite realizability for simple graphs when degrees are bounded only by n1n-1 [barnoy2]. Our results contribute to this open direction by showing that if bipartite realizability for simple graphs with degrees bounded by n1n-1 were NP-complete, then NP-completeness would persist even under much stronger restrictions, namely for any upper bound of the form $cn$ for any fixed constant $c>0$.
@@ 2 Preliminaries
We introduce the necessary ingredients to achieve the stated results. We are going to change graphs using hinge-flips. A hinge-flip operation removes an edge $(u_{1},v)$ and add an edge $(u_{2},v)$ not presented in the edge set before the operation. The hinge-flip operations do not change the sum of the degrees, however, they do change the degree sequence of the graph on which they act. However, observe the following: if in a pair of degree sequences $(D_{1},D_{2})$ the degree of $u_{i}$ is greater than the degree of $u_{j}$, then in any realization of $(D_{1},D_{2})$ there exists a hinge-flip that removes an edge $(u_{i},v)$ and adds an edge $(u_{j},v)$. That is, if a pair of degree sequences $(D_{1},D_{2})$ is bigraphic than any pair of degree sequences $(D_{1}^{\prime},D_{2})$ is also bigraphic that can be obtained from $(D_{1},D_{2})$ by decreasing a degree $d_{j}\in D_{1}$ and increasing a $d_{i}\in D_{1}$ with $d_{i}<d_{j}$. When $d_{i}<d_{j}-1$, we call the corresponding hinge-flip a balancing hinge-flip (if $d_{i}=d_{j}+1$, the hinge flip only swap the two degrees). Observe that a balancing hinge-flip does not change the sum of the degrees, cannot decrease the minimum degree as well as cannot increase the maximum degree. This observation leads to the notation of the least balanced degree sequence and a lemma on it.
Definition 2.1.
@@
Lemma 2.2.
Let $(D_{1},D_{2})$ be a pair of degree sequences. If $D_{1}\in\mathcal{D}(n_{1},\Sigma,\Delta_{1},d_{1})$, $D_{2}\in\mathcal{D}(n_{2},\Sigma,\Delta_{2},d_{2})$ and $(LBDS(n_{1},\Sigma,\Delta_{1},d_{1}),LBDS(n_{2},\Sigma,\Delta_{2},d_{2}))$ is bigraphic then $(D_{1},D_{2})$ is also bigraphic.
The proof can be found in [limiklos]; an analogous observation on simple degree sequences was already observed by Aigner and Triesch [at94].
The graphicality of a bipartite degree sequence can be checked with the Gale-Ryser theorem [gale, ryser]. The Gale-Ryser theorem says that a bipartite degree sequence $D:=(D_{1},D_{2})$ with $D_{1}=d_{1,1}\geq d_{1,2}\geq\ldots\geq d_{1,n}$ and $D_{2}=d_{2,1},d_{2,2},\ldots,d_{2,m}$ is bigraphic if and only if the sum of the degrees in $D_{1}$ equals the sum of the degrees in $D_{2}$ and for all $k=1,2,\ldots,n$,
(1) $\sum_{i=1}^{k}d_{1,i}\leq\sum_{j=1}^{m}\min\{k,d_{2,j}\}.$
We make the following observation. A bipartite degree sequence $D:=(D_{1},D_{2})$ with $D_{1}=d_{1,1}\geq d_{1,2}\geq\ldots\geq d_{1,n}$ and $D_{2}=d_{2,1},d_{2,2},\ldots,d_{2,m}$ is bigraphic if and only if the degree sequence $D^{\prime}:=d_{1,1}+n-1,d_{1,2}+n-1,\ldots,d_{1,n}+n-1,d_{2,1},d_{2,2},\ldots,d_{2,m}:=$ $d^{\prime}_{1},d^{\prime}_{2},\ldots d^{\prime}_{n^{\prime}}$ is graphic. Then for all $k\leq n$, the Gale-Ryser inequalities in equation 1 are equivalent with the Erdős-Gallai inequalities [eg]:
(2) $\sum_{i=1}^{k}d^{\prime}_{i}\leq k(k-1)+\sum_{j=k+1}^{n^{\prime}}\min\{k,d_{j}\}.$
@@
Tripathi and Vijay [TV03] proved that a degree sequence is graphic if and only if the sum of the degrees is even and for all $k$ with $d_{k}>d_{k+1}$ the Erdős-Gallai inequalities hold. It leads to the following observation. {observation}
The bipartite degree sequence
$D=(LDBS(n_{1},\Sigma,\Delta_{1},d_{1}),LDBS(n_{2},\Sigma,\Delta_{2},d_{2}))$
is bigraphic if and only if the Gale-Ryser inequality with $k=\left\lfloor\frac{\Sigma-n_{1}d_{1}}{\Delta_{1}-d_{1}}\right\rfloor$ and $k=\left\lceil\frac{\Sigma-n_{1}d_{1}}{\Delta_{1}-d_{1}}\right\rceil$ are satisfied.
@@ 3 Polynomial algorithm for $\textsc{BDR}_{c_{1},c_{2}}$
We are going to prove the tractable result of the parameterized BDR problem. Before we state and prove the theorem on the tractable result, we need a technical observation. {observation} Let $c_{1}\geq 0$ and $0\leq c_{2}<1$. Then
$c_{1}\;\geq\;\frac{c_{2}^{2}}{1-c_{2}}\quad\text{if and only if}\quad c_{2}\;\leq\;\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}.$
Proof 3.1.
Since $0\leq c_{2}<1$, the denominator $1-c_{2}$ is positive. Therefore,
$c_{1}\geq\frac{c_{2}^{2}}{1-c_{2}}$
is equivalent to
(3) $c_{1}(1-c_{2})\geq c_{2}^{2}.$
Rearranging (3), we obtain
$c_{2}^{2}+c_{1}c_{2}-c_{1}\leq 0.$
We now view this as a quadratic inequality in the variable $c_{2}$. The corresponding quadratic equation
$c_{2}^{2}+c_{1}c_{2}-c_{1}=0$
has discriminant
$\Delta=c_{1}^{2}+4c_{1}=c_{1}(c_{1}+4).$
Thus, the two real roots are
$c_{2}=\frac{-c_{1}\pm\sqrt{c_{1}(c_{1}+4)}}{2}.$
Since $c_{2}\geq 0$, only the positive root is relevant. Therefore,
$c_{2}^{2}+c_{1}c_{2}-c_{1}\leq 0\quad\Longleftrightarrow\quad 0\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}.$
This establishes the claimed equivalence.
Theorem 3.2.
@@
Observe the following. The inequality $c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$ implies that $c_{1}\leq\frac{1}{2}$ because for any $c_{1}>\frac{1}{2}$ it holds that $c_{1}>\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$. When $0\leq c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$ , it is enough to show that for any $s_{1}\leq s_{2}$ with $s_{1}+s_{2}=n$, and for any $\Sigma$ with $c_{1}s_{2}\leq\Sigma\leq c_{2}s_{1}$, the bipartite degree sequence $D:=(LBDS(s_{1},\Sigma,c_{1}n,c_{2}n),LBDS(s_{2},\Sigma,c_{1}n,c_{2}n))$ is graphic. Indeed, according to Lemma 2.2, then any degree sequence $D^{\prime}:=(D_{1},D_{2})$ with $D_{1}\in\mathcal{D}(s_{1},\Sigma,c_{1}n,c_{2}n)$ and $D_{2}\in\mathcal{D}(s_{2},\Sigma,c_{1}n,c_{2}n)$ is graphic. Therefore, a degree sequence $D\in\mathcal{D}(n,2\Sigma,c_{1}n,c_{2}n)$ has a bipartite graph realization if and only if $D$ has a split of two degree sequences with equal sum. This is the subset sum problem, which is NP-complete in general, but has a pseudo-polynomial time algorithm. The pseudopolinomiality means that the runtime of the algorithm grows polynomial with the values of the numbers and not the number of digits needed to describe the numbers. However, any degree is at most $\frac{n}{2}$, therefore the subset sum problem is actually in P for these degree sequences.
We use the Gale-Ryser theorem to prove that the bipartite degree sequence $D$ is graphic. According to Observation 2, it is enough to check the Gale-Ryser inequalities with $k=\left\lfloor\frac{\Sigma-s_{1}c_{1}n}{(c_{2}-c_{1})n}\right\rfloor$ and $k=\left\lceil\frac{\Sigma-s_{1}c_{1}n}{(c_{2}-c_{1})}\right\rceil$.
@@
$c_{1}\geq\frac{c_{2}^{2}}{1-c_{2}},$
which is equivalent with $c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$, according to Observation 3. There are possible intermediate degrees with
@@
We have to check those Gale-Ryser inequalities which were obtained in Observation 2. The first is:
(5) $n_{1}c_{2}n\;\leq\;n_{2}\min\{n_{1},c_{2}n\}\;+\;\min\{n_{1},d_{2}\}\;+\;m_{2}\min\{n_{1},c_{1}n\}.$
We now prove equation 5 by splitting into three exhaustive cases according to the relative size of $n_{1}$.
@@
In this range we have $\min\{n_{1},c_{2}n\}=c_{2}n$, $\min\{n_{1},d_{2}\}=d_{2}$ (because $d_{2}<c_{2}n<n_{1}$) and $\min\{n_{1},c_{1}n\}=c_{1}n$. Thus, the right-hand side of equation 5 equals
@@
This is exactly the desired inequality in Case (A). Hence equation 5 holds when $n_{1}\geq c_{2}n$.
@@
Here $\min\{n_{1},c_{2}n\}=n_{1}$, $\min\{n_{1},d_{2}\}=n_{1}$ (since $d_{2}\geq c_{1}n\geq n_{1}$), and $\min\{n_{1},c_{1}n\}=n_{1}$. Therefore the right-hand side of equation 5 equals
@@
However, we know that $c_{2}\leq\frac{c_{1}}{c_{2}+c_{1}}$ for it is equivalent with the condition $\frac{c_{2}^{2}}{1-c_{2}}\leq c_{1}$. Therefore, equation 5 holds in Case (B).
@@
In this range we have $\min\{n_{1},c_{2}n\}=n_{1}$, and $\min\{n_{1},c_{1}n\}=c_{1}n$. Hence equation 5 becomes
@@
(7) $n_{1}c_{2}n+d_{1}+\left(\frac{n}{2}-n_{1}+x-1\right)c_{1}n=(n_{1}+y)c_{2}n+d_{2}+\left(\frac{n}{2}-n_{1}-x-y-1\right)c_{1}n.$
@@
$n_{1}c_{2}n\leq\left(n_{1}+\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}\right)n_{1}+\min\{n_{1},d_{2}\}+$
$\left(\frac{n}{2}-n_{1}-x-\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}-1\right)c_{1}n.$
@@
$0$ $\leq$ $\underbrace{\eta^{2}-(c_{1}+c_{2})\eta+\frac{c_{1}}{2}}_{R_{1}}+\underbrace{\frac{\frac{d_{1}-d_{2}}{n^{2}}}{c_{2}-c_{1}}(\eta-c_{1})+\frac{\min\{\eta,\frac{d_{2}}{n}\}}{n}}_{R_{2}}+$
$+\underbrace{z\left(\frac{2c_{1}}{c_{2}-c_{1}}\eta-\left(1+\frac{2c_{1}}{c_{2}-c_{1}}\right)c_{1}\right)}_{R_{3}}.$
@@
Therefore $R_{2}$ is always positive. Therefore, we can look at equation 3.3 as